OpenStudy (anonymous):

What is the solution to the equation 36^3c = square root of 6^(8c – 6) ?

mathslover (mathslover):

\[\large{36^{3c}=\sqrt{6^{8c-6}}}\] is this the question?

OpenStudy (anonymous):

6 sqrt (8c - 6)

mathslover (mathslover):

oh k : \[\large{36^{3c}=6\sqrt{8c-6}}\] like this?

OpenStudy (anonymous):

Yes

OpenStudy (lgbasallote):

\[\huge 36^{3c} = 6^{\sqrt{8c-6}}\] sounds so much cooler

OpenStudy (anonymous):

OK

mathslover (mathslover):

6c = sqrt{8c-6}

OpenStudy (anonymous):

Can you do more to it?

mathslover (mathslover):

yes wait

mathslover (mathslover):

\[\large{36 c^2 = 8c-6}\] \[\large{36c^2 - 8c+6=0}\]

OpenStudy (lgbasallote):

woah...what happened there...

OpenStudy (anonymous):

?

mathslover (mathslover):

r u confused? where?

OpenStudy (anonymous):

c= -3/2